OpenStudy (anonymous):

Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of 14 inches and a height of 4 inches. Each leg is 4 inches by 4 inches by 18 inches. how many area will Evan have to paint?

OpenStudy (anonymous):

are you looking for just an answer or explanation too?

OpenStudy (anonymous):

both .. if thats okay

OpenStudy (anonymous):

okay just wait for a while explaining may take a bit :)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Total surface area of cylinder= \[2 \pi r ^{2} + 2 \pi r h \]= 483.80 sqinch total surface area for legs= 4[2(4x4)+4(4x18)]=1280 sqinch Now subtract the overlapping area between joining of legs onto the top. . . . overlapping = 8(4x4)= 128 sq inch So area to be painted= 483.80+1280-128= 1635.8

OpenStudy (anonymous):

thanks!